# Is it Really Negative? Evaluating Natural Language Video Localization Performance on Multiple Reliable Videos Pool

### v2

@@ Abstract
With the explosion of multimedia content in recent years, Video Corpus Moment Retrieval (VCMR), which aims to detect a video moment that matches a given natural language query from multiple videos, has become a critical problem. However, existing VCMR studies have a significant limitation since they have regarded all videos not paired with a specific query as negative, neglecting the possibility of including false negatives when constructing the negative video set. In this paper, we propose an MVMR (Massive Videos Moment Retrieval) task that aims to localize video frames within a massive video set, mitigating the possibility of falsely distinguishing positive and negative videos. For this task, we suggest an automatic dataset construction framework by employing textual and visual semantic matching evaluation methods on the existing video moment search datasets and introduce three MVMR datasets. To solve MVMR task, we further propose a strong method, CroCs, which employs cross-directional contrastive learning that selectively identifies the reliable and informative negatives, enhancing the robustness of a model on MVMR task. Experimental results on the introduced datasets reveal that existing video moment search models are easily distracted by negative video frames, whereas our model shows significant performance. [^1] We have opened an anonymous GitHub page to make our code and dataset publicly available: link
1 Introduction
@@
Enabled by the increased accessibility of video-sharing platforms along with the advancement of networking and storage technology, a vast number of new content is available on the web daily. With this huge number of contents, Natural Language Video Localization (NLVL) task, searching for a video moment suitable for a given query, has emerged as an essential problem (Anne Hendricks et al., 2017; Gao et al., 2017). However, this line of research narrows down the task to an impractical setting by targeting the search over a single video; thus, it bears limitations in retrieving relevant information over whole media pools.
Figure 1: MVMR vs. VCMR. In existing VCMR studies, only a single golden positive video moment is identified as positive, while entire videos not paired with a specific query are regarded as negative. Our proposed MVMR filters positive video moments for a query from the whole video set; thus, it covers more practical and generalized use cases with reliable settings. $v^{+}_{i}$ and $v^{-}_{j}$ mean a positive and a negative video, respectively.
Therefore, Video Corpus Moment Retrieval (VCMR) task that retrieves a correct positive moment among numerous negative videos has emerged (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022). Nonetheless, VCMR also has significant limitations as it merely extends NLVL datasets by categorizing all unpaired moments for a specific text query as negative. This assumption diverges from the practical setting of video moment search where multiple positive videos exist in whole media collections. Furthermore, existing VCMR studies have overlooked the potential presence of false-negative videos within the unpaired video set when constructing video retrieval pools for a specific query. As a result, evaluating the performance of NLVL models using an existing VCMR framework may lead to unreliable results. For example, in TACoS test dataset, it is noteworthy that the number of videos with queries exactly matching the text "The person gets out a knife." amounts to 60% of all the videos. It is a significant problem since more videos will likely be incorrectly classified as negatives when expanding the matching criteria to include broader semantic similarities.
@@
To bridge these gaps, we propose an MVMR (Massive Videos Moment Retrieval) task, which expands the search coverage to a massive video set that can include any number of positive videos as shown in Figure 1. Specifically, given a query $q$ with a positive videos $v^{+}_{1},...,v^{+}_{k}$ and a negative videos $v^{-}_{1},...,v^{-}_{n-k}$, the MVMR task aims to detect temporal moments in the positive videos that matches the query from a massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$. Furthermore, we propose simple and effective methods to construct a practical MVMR dataset by leveraging publicly available NLVL datasets. A critical challenge in converting an NLVL dataset into an MVMR dataset is accurately defining positive and negative videos for each query. However, manually performing this task by humans is exceedingly labor-intensive due to the massive volume of videos that need to be evaluated for each query. To address this issue, we propose an automated framework for MVMR dataset construction. This framework employs textual and visual semantic matching evaluation methods to define MVMR positive and negative videos while mitigating the danger of false categorization. Specifically, we examine the similarity between a target query and all videos in an NLVL dataset using SimCSE Gao et al. (2021) and EMScore Shi et al. (2022) models to define a massive video set. To this end, we construct three MVMR datasets, and we confirm that three samples of the introduced datasets contain only 1.5%, 0.2%, and 3.5% falsely defined videos, showing that our approach successfully creates a practical benchmark dataset from human evaluation.
To tackle MVMR task, we also introduce a novel negative sample training method, Cross-directional Hard and Reliable Negative Contrastive Learning (CroCs). Specifically, CroCs adopts two training mechanisms: (1) weakly-supervised potential negative learning by assessing the similarity between queries, and (2) hard-negative learning by utilizing the relevance score between queries and moments predicted by a model, which helps find challenging negative samples. We demonstrate the performance of strong baseline models, including a state-of-the-art model, to compare with our model in the new scenario (MVMR). The experimental results show that the performance of baseline models significantly degrades on the new task. For example, the state-of-the-art model, MMN, shows average 29.7% degradation on R@1(IOU0.5) for all MVMR datasets, revealing the difficulty of the MVMR task. In contrast, our CroCs outperforms baseline models significantly in the MVMR setting, and this result proves that informative negative samples enhance the performance of the MVMR task. Moreover, we solve the MVMR task in a more practical scenario by adopting a pipeline with a video retrieval model, revealing that CroCs also outperforms the strong baseline.
2 Backgrounds
2.1 Natural Language Video Localization
@@
Natural language video localization (NLVL), which aims to detect a specific moment in a video that matches a given text query, is one of the most challenging problems in the video-language multi-modal domain (Zhang et al., 2020b, a; Nan et al., 2021; Gao and Xu, 2021; Liu et al., 2021; Wang et al., 2022). Therefore, the NLVL task has been explored extensively, with numerous effective methods emerging. Some novel approaches have regarded the problem as question-answering (QA) tasks (Seo et al., 2018; Joshi et al., 2020), by leveraging a QA model to encode a multi-modal representation and then predicting the frames that correspond to the start and end of the moment that match a query (Ghosh et al., 2019; Zhang et al., 2020a). However, NLVL studies operate within an impractical framework, focusing on searching within a single video. Consequently, they have limitations in retrieving relevant information across the whole media repositories.
@@
Table 1: Summary of datasets. Remarkably, the queries of $\text{MVMR}_{\text{Charades-STA}}$ and $\text{MVMR}_{\text{TACoS}}$ generally include multiple positive moments, and it proves that existing VCMR studies are exposed to the risk of false negatives.
2.2 Video Corpus Moment Retrieval
Existing studies (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022) have tried to extend the NLVL to search on multiple video pools, naming it the Video Corpus Moment Retrieval (VCMR) task, which focuses on retrieving a moment from multiple videos for a given query. However, the VCMR simply extends the NLVL task by designating only one matched positive video moment for a target query as a positive, while labeling all unmatched videos as negative videos. This approach deviates from the practical scenario of the video moment search since multiple positive videos can be present for a target query. Moreover, existing VCMR studies have failed to consider the possibility of false-negative when constructing the negative video set for a specific query, which poses a significant reliability issue.
3 Problem Definition
MVMR task evaluates video moment retrieval models under an MVMR dataset with $k$ positive videos and $n-k$ negative videos for each query. Given a positive video set $V^{+}_{q}=\{v^{+}_{1},...,v^{+}_{k}\}$ and a negative video set $V^{-}_{q}=\{v^{-}_{1},...,v^{-}_{n-k}\}$, the MVMR task aims to localize a moment $(x_{s},x_{e})$ of a specific video $v$ that matches the query from massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$, where $x_{s}$ and $x_{e}$ mean start and end points of the moment, respectively. In MVMR task, we aim to retrieve an optimal positive moment $(x^{*}_{s},x^{*}_{e})$ that matches semantically with the text query $q$. We derive confidence scores of moments for the whole $n$ massive videos and select the moment with the highest score as a prediction.
4 MVMR Dataset Construction
@@
Our MVMR setting automatically reconstructs an NLVL dataset by expanding it to associate each query with multiple positive and negative moments. The paired positive and negative videos are carefully selected with semantic similarity check techniques. In the following, we detail our process for constructing an MVMR dataset and present the analyses of the three constructed MVMR datasets.
4.1 Semantic Similarity Check
@@
This section examines the similarity between a target query $q$ and a video $v$ using semantic text embedding and video captioning evaluation models.
Query Similarity-based Check.
@@
NLVL datasets consist of video set $V=\{v_{1},...,v_{l}\}$ and paired query set $Q_{v}=\{q_{1},...,q_{m}\}$ for each video $v$, where $l$ and $m$ is the number of videos and text queries included in a dataset, respectively. Since each text query semantically describes the paired moment, we use queries to obtain semantic information about the paired video. For example, suppose that a video $v$ is paired with a query $q$, "The person gets out a knife", in a moment $(x_{s},x_{e})$. Since a target query $\hat{q}$, "The person takes out a knife", is highly similar to the query $q$, $\hat{q}$ is associated with the moment $(x_{s},x_{e})$ of the video $v$. We use a semantic text embedding model, SimCSE Gao et al. (2021), to distinguish positive and negative videos for each query. Specifically, we first obtain embeddings $e_{\hat{q}}$ and $e_{q_{1}},...,e_{q_{m}}$ for the target text query $\hat{q}$ and all video text queries $q_{1},...,q_{m}$ using SimCSE, respectively. We calculate cosine similarity $s_{te}(\hat{q},q_{i})$ between $e_{\hat{q}}$ and $e_{q_{i}}$ to quantify the semantic matching, and select the maximum similarity among $s_{te}(\hat{q},q_{1}),...,s_{te}(\hat{q},q_{m})$ to use it as the query-video matching score $s_{te}(\hat{q},v)$ of the target query $\hat{q}$ and the video $v$. We select max aggregation to obtain $s_{te}(\hat{q},v)$ since the video should be regarded as a positive video, even if only one similar query to a target query exists in a video.
Figure 2: Examples of constructed MVMR datasets. We visualize positive and negative video sets for the query “person closes the door" of the $\text{MVMR}_{\text{Charades-STA}}$. A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive. The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
Query-video Similarity-based Check.
@@
Queries associated with a video may describe the video limitedly since they may not cover all of its context. Therefore, we examine a visual semantic matching between a query and a video to filter positive and negative videos. We use EMScore Shi et al. (2022), a video captioning evaluation model, to quantify the visual semantic similarity $s_{tv}$, used as a complementary filtering method on videos primarily filtered through the query similarity $s_{te}$.
4.2 Distinguishing Positive and Negative
@@
This section describes constructing positive and negative video sets using the calculated query-video similarity, $s_{te}$ and $s_{tv}$. Positive candidates are distinguished by primarily excluding videos $s_{te}(q,v)<t_{te}^{+}$, where $t_{te}^{+}$ is a hyper-parameter. We further exclude videos that have lower similarity than the mean $s_{tv}$ of all golden positive query-video pairs $(q,v^{*})$ for all queries. We construct final MVMR positives by randomly sampling $k$ videos from this filtered set of positive candidates. Negative candidates are defined by excluding videos with $s_{te}(q,v)>t_{te}^{-}$, where $t_{te}^{-}$ is also a hyper-parameter. We additionally exclude videos that have higher similarity than the mean $s_{tv}$ of query-video pairs $(q,v^{-})$, which are primarily filtered as negatives using $s_{te}$. Like MVMR positives, we construct final MVMR negatives by randomly sampling $(n-k)$ videos from this set of negative candidates. From this process, a total of $n$ videos retrieval pool $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$ is obtained for each query.
5 MVMR Dataset Analysis
5.1 MVMR Settings.

### v4

# MVMR: A New Framework for Evaluating Faithfulness of Video Moment Retrieval against Multiple Distractors

@@ Abstract.
With the explosion of multimedia content, video moment retrieval (VMR), which aims to detect a video moment that matches a given text query from a video, has been studied intensively as a critical problem. However, the existing VMR framework evaluates video moment retrieval performance, assuming that a video is given, which may not reveal whether the models exhibit overconfidence in the falsely given video. In this paper, we propose the MVMR (Massive Videos Moment Retrieval for Faithfulness Evaluation) task that aims to retrieve video moments within a massive video set, including multiple distractors, to evaluate the faithfulness of VMR models. For this task, we suggest an automated massive video pool construction framework to categorize negative (distractors) and positive (false-negative) video sets using textual and visual semantic distance verification methods. We extend existing VMR datasets using these methods and newly construct three practical MVMR datasets. To solve the task, we further propose a strong informative sample-weighted learning method, CroCs, which employs two contrastive learning mechanisms: (1) weakly-supervised potential negative learning and (2) cross-directional hard-negative learning. Experimental results on the MVMR datasets reveal that existing VMR models are easily distracted by the misinformation (distractors), whereas our model shows significantly robust performance, demonstrating that CroCs is essential to distinguishing positive moments against distractors. Our code and datasets are publicly available: link.
Figure 1. Faithfulness Issue in Video Moment Retrieval. If a VMR model is trained without considering overconfidence in a given context (video), it may undergo a Type 1 error since it prioritizes selecting a video moment from the falsely given video (distractor).
Video Moment Retrieval, Trustworthy AI, Contrastive Learning
submissionid: 2163journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679838ccs: Information systems Multimedia and multimodal retrieval
1. Introduction
@@
Enabled by the increased accessibility of video-sharing platforms along with the advancement of networking and storage technology, a vast amount of new content becomes available on the web daily. With this huge number of contents, Video Moment Retrieval (VMR) task, searching for a video moment suitable for a given query, has emerged as an essential problem (Anne Hendricks et al., 2017; Gao et al., 2017).
However, this line of research evaluates retrieval models assuming that a context (video) is given, it is limited in revealing the capability of distinguishing misinformation. Figure 1 exemplifies the problem of faithfulness for misinformation demonstrated by a VMR model. This overconfidence issue has been studied recently in the natural language community as a challenging task (Chen et al., 2022; Shi et al., 2023), but it has not been studied sufficiently in the VMR field. This problem also arises in a practical video moment retrieval scenario, Video Corpus Moment Retrieval (VCMR), which aims to search for a relevant video moment within a pool of multiple videos (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022). The VCMR framework evaluates video moment retrieval performance using a sequential pipeline that integrates video retrieval with video moment retrieval. Therefore, it also diverges from evaluating the overconfidence of a VMR model, as the evaluated retrieval performance may predominantly depend on video retrieval capability.
Figure 2. Massive Video Pool Construction. From our experiment on the TACoS dataset, if a massive video set is constructed using random sampling, then over 40% of queries include at least one false-negative video, showing the risk of random sampling. Our method constructs a massive video set by filtering trustful positive and negative sets using semantic distance verification methods, considering the possibility of false-negatives. $v^{+}_{i}$ and $v^{-}_{j}$ mean a positive and a negative video, respectively.
To tackle this issue, we propose an MVMR (Massive Videos Moment Retrieval for Faithfulness Evaluation) task, which expands the search coverage to a massive video set that can include any number of positive and negative videos for evaluating the faithfulness of VMR models. Specifically, given a text query $q$ with positive videos $v^{+}_{1},...,v^{+}_{k}$ and negative videos $v^{-}_{1},...,v^{-}_{n-k}$, the MVMR task aims to retrieve temporal moments in the positive videos that match the query from a massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$ without the interference of any video retrieval model.
@@
For the MVMR task, we further construct practical MVMR benchmarks by extending publicly available VMR datasets. Specifically, we categorize positive and negative videos for a specific query from the whole video set of a VMR dataset to construct a massive video set. The most critical challenge in converting a VMR dataset into an MVMR dataset is accurately categorizing positive and negative videos for each query. However, addressing this issue by manually annotating a large number of videos for each query to determine whether they are positive or negative is exceedingly labor-intensive. Therefore, we propose a new automated framework to construct MVMR datasets by employing textual and visual semantic distance verification methods to categorize positive and negative videos for each query, as shown in Figure 2. Specifically, we examine the distance between a target query and all videos in a VMR dataset using state-of-the-art text embedding model (Gao et al., 2021) and video-captioning evaluation model (Shi et al., 2022) to define the positive and negative massive video sets. Consequently, we construct three MVMR datasets, and we confirm that the introduced datasets contain only 1.5%, 0.2%, and 3.5% falsely defined videos from human evaluation, showing that our approach successfully constructs practical benchmarks. On these benchmarks, we evaluate strong baseline VMR models, including state-of-the-art models, and demonstrate that significant overconfidence issues arise in the new scenario (MVMR). For example, the state-of-the-art model, MMN, shows an average 35.9% degradation in the performance (R1@0.5 and R5@0.5) for all MVMR datasets, revealing the difficulty of the MVMR task.
In addition, to enhance the faithfulness of VMR models, we introduce a novel informative sample-weighted learning, CroCs, which stands for Cross-directional informative sample-weighted Contrastive Learning. Specifically, CroCs adopts two training mechanisms: (1) weakly-supervised potential negative learning and (2) cross-directional hard-negative learning. The weakly-supervised potential negative learning assesses the distance between queries to exclude false-negatives. The cross-directional hard-negative learning measures the relevance score between queries and moments predicted by a model to exclude easy negative samples.
We demonstrate that our CroCs outperforms strong baselines significantly in the MVMR setting, and this result proves that informative sample-weighted constrastive learning successfully mitigates the misinformation overconfidence of a VMR model. Moreover, we show that CroCs is also effective in the realistic scenario by adopting the widely-used pipeline, which includes a video retrieval model. In summary, this work makes the following contributions:
  • We propose a new task called MVMR, and construct three MVMR datasets to evaluate the faithfulness of VMR models.
  • We propose a novel automated dataset construction framework, the first to consider how to construct a trustful massive video retrieval pool via semantic distance verification.
  • We demonstrate that existing VMR models are significantly vulnerable to misinformation; thus, we introduce a novel informative sample-weighted learning method, CroCs, to enhance the faithfulness of VMR models against misinformation.
Figure 3. Examples of constructed MVMR datasets. We visualize positive and negative video sets for a query of the constructed three MVMR datasets. A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive. The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
2. Video Moment Retrieval
@@
Video Moment Retrieval (VMR), which aims to detect a specific moment in a video that matches a given text query, is one of the most challenging problems in the video-language multi-modal domain. Therefore, the VMR task has been explored extensively using numerous effective methods  (Zhang et al., 2020a, b; Nan et al., 2021; Gao and Xu, 2021; Liu et al., 2021; Wang et al., 2022; Li et al., 2023). Some novel approaches have regarded the problem as question-answering (QA) tasks (Seo et al., 2018; Joshi et al., 2020), by leveraging a QA model to encode a multi-modal representation and then predicting the frames that correspond to the start and end of the moment that match a query (Ghosh et al., 2019; Zhang et al., 2020b). However, VMR studies operate within an impractical framework, focusing on searching moments within a given video. Consequently, their settings are limited in retrieving relevant information across all media repositories and have not evaluated the faithfulness of VMR models on misinformation.
Existing studies (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022) have extended the VMR to search on a multiple video pool, naming the task as the Video Corpus Moment Retrieval (VCMR), which focuses on retrieving a moment from multiple videos for a given query. The VCMR task considers the most practical usage scenario of VMR. However, it is still limited by its inability to evaluate the faithfulness of VMR models in a misinformative context. Existing VCMR studies have utilized a framework evaluating moment retrieval performance by integrating video retrieval with VMR; thus, the evaluated performance in the VCMR setting may overwhelmingly depend on video retrieval capability. In addition, the VCMR framework simply extends the VMR task by categorizing only a paired positive video moment for a target query as positive while labeling all unpaired videos as negative videos (Yoon et al., 2023; Zhang et al., 2021). This approach deviates from the practical scenario of the video moment search since multiple positive videos can be present in whole media collections for a target query. Moreover, existing VCMR studies have randomly selected negative samples from all unpaired videos; thus, they have overlooked the possibility of false-negatives when constructing the negative video set for a query, which raises significant trustfulness concerns.
Some studies (Yoon et al., 2022; Hao et al., 2022; Yoon et al., 2023) have attempted to address unusual behaviors of VMR models, such as bias or spurious correlation, which arise during the training for the VMR and VCMR tasks. Yoon et al. (2022) has categorized good and bad biases to train a model using only a good bias. Hao et al. (2022) has mitigated a temporal bias by training a model with temporally shuffled moments. However, existing studies have never addressed the overconfidence issue in a misinformative context; thus, it remains a challenging task to solve.
3. MVMR Problem Definition
The MVMR task evaluates VMR models under a massive video pool with $k$ positive videos and $n-k$ negative videos for each query $q$. Given a positive video set $V^{+}_{q}=\{v^{+}_{1},...,v^{+}_{k}\}$ and a negative video set $V^{-}_{q}=\{v^{-}_{1},...,v^{-}_{n-k}\}$, the MVMR task aims to search for a moment $m=(x_{s},x_{e})$ of a specific video $v$ that matches the query from massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$, where $x_{s}$ and $x_{e}$ mean start and end points of the moment, respectively. In contrast to the VCMR task, we only utilize a VMR model $\mathcal{P}$ by inputting a target query $q$ and each video $v$ in $V^{+,-}_{q}$ to compute the matching score $p=\mathcal{P}(m|q,v)$ for each moment $m$. We finally determine the optimal moment $m^{*}=(x^{*}_{s},x^{*}_{e})$ by comparing the matching score of all moments in $V^{+,-}_{q}$ as $m^{*}=argmax_{m}\mathcal{P}(m|q,v)$.
@@ 3. MVMR Problem Definition
Table 1. Summary of datasets. Remarkably, the queries of $\text{MVMR}_{\text{Charades-STA}}$ and $\text{MVMR}_{\text{TACoS}}$ generally include multiple positive moments, and it empirically proves that a massive video set construction using random sampling is inadequate to construct the MVMR dataset since it is exposed to the risk of false negatives.
4. MVMR Dataset Construction
@@
Our MVMR setting automatically reconstructs a VMR dataset by expanding it to associate each query with multiple positive and negative moments. Each VMR dataset consists of multiple videos with some query-moment pairs, and our method carefully categorizes the positive and negative sets using two semantic distance verification techniques: Query Distance-based Verification and Query-video Distance-based Verification. In the following, we detail our process for constructing the MVMR benchmark and present the analyses of three constructed MVMR datasets.
4.1. Semantic Distance Verification
@@
This section describes the verification process of the distance between a target query $q$ and a video $v$ using semantic text embedding and video captioning evaluation models.
Query Distance-based Verification.
@@
VMR datasets consist of video set $V=\{v_{1},...,v_{l}\}$ and paired query set $Q_{v}=\{q_{1},...,q_{m}\}$ for each video $v$, where $l$ and $m$ is the number of videos and text queries included in a dataset, respectively. Since each text query semantically describes the paired moment, we use queries to obtain semantic information about the paired video. For example, suppose that a video $v$ is paired with a query $q$, ”The person gets out a knife”, in a moment $(x_{s},x_{e})$. Since a target query $\hat{q}$, ”The person takes out a knife”, is highly similar to the query $q$, $\hat{q}$ is associated with the moment $(x_{s},x_{e})$ of the video $v$. We use a semantic text embedding model, SimCSE (Gao et al., 2021), to distinguish positive and negative videos for each query. Specifically, we first obtain embeddings $e_{\hat{q}}$ and $\{e_{q_{1}},...,e_{q_{m}}\}$ for the target text query $\hat{q}$ and all video text queries $\{q_{1},...,q_{m}\}$ using SimCSE, respectively. Then, we calculate cosine similarity $s_{te}(\hat{q},q_{i})$ between $e_{\hat{q}}$ and $e_{q_{i}}$ to quantify the semantic matching, and select the maximum similarity among $\{s_{te}(\hat{q},q_{1}),...,s_{te}(\hat{q},q_{m})\}$ to use it as the query-video matching score $s_{te}(\hat{q},v)$ of the target query $\hat{q}$ and the video $v$. We select max aggregation to obtain $s_{te}(\hat{q},v)$ since the video should be regarded as a positive video, even if only one similar query to a target query exists in a video.
Query-video Distance-based Verification.
@@
Paired queries may describe the video limitedly since they may not cover all its context. We examine a visual semantic matching between a query and a video to categorize positive and negative videos. We use EMScore (Shi et al., 2022), a video captioning evaluation model, to quantify the visual similarity $s_{tv}(q,v)$, used as a complementary filtering method on videos primarily selected through the query similarity $s_{te}(q,v)$.
4.2. Categorizing Positive and Negative sets
@@
This section describes constructing positive and negative video sets using the calculated query-video similarities, $s_{te}$ and $s_{tv}$. Positive candidates are identified by primarily excluding videos with $s_{te}(q,v)<t_{te}^{+}$, where $t_{te}^{+}$ is a hyper-parameter. We further exclude videos that have lower similarity than the mean $s_{tv}$ of all golden positive query-video pairs $(q,v^{*})$ for all queries. We construct final MVMR positives by randomly sampling $k$ videos from this categorized set of positive candidates. Negative candidates are identified by excluding videos with $s_{te}(q,v)>t_{te}^{-}$, where $t_{te}^{-}$ is also a hyper-parameter. We additionally exclude videos that have higher similarity than the mean $s_{tv}$ of query-video pairs $(q,v^{-})$, which are primarily categorized as negatives using $s_{te}$. Like MVMR positives, we construct final MVMR negatives by randomly sampling $(n-k)$ videos from this set of negative candidates. From this process, a total of $n$ video retrieval pool $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$ is obtained for each query.
5. MVMR Dataset Analysis
5.1. MVMR Settings.